What volume (L) of O2 at 760. mmHg and 28 0C is needed to react fully with 3.8 g of C2H6 (ethane) at the same temperature and pressure?
1
Expert's answer
2020-04-27T01:45:35-0400

Convert mass of C2H6 to moles

n(C2H6) = m/M = 3.8g/30g/mol = 0.1267 mol

2C2H6 + 7O2 = 4CO2 + 6H2O

According to the equation, use simple proportion to find moles O2 required

2 mol C2H6 - 7 mol O2

0.1267 mol - x

n(O2) = 0.443 mol

Convert moles of O2 to a volume

V(O2) = n*Vm = 0.443mol*22.4L/mol = 9.92 L